OpenStudy (anonymous):

reward!!!!!!!!!!!! just help give me a medal too if u want one Twelve and one-fourth divided by two and one-tenth equals what number? five and five-sixths six and one-fortieth six and two-fifths six and one-fourth

OpenStudy (anonymous):

Twelve and one-fourth divided by two and one-tenth equals what number? 12 1/4 / 2 1/10 =49/4 / 21/10 =245/20 / 42/20 = 245/42 = 35/6 = 5 5/6 Answer: five and five-sixths

OpenStudy (anonymous):

A